# Resolving Availability and Run-time Integrity Conflicts in Real-Time Embedded Systems

### v1

@@ 6. Implementation & Verification
PAIR modules are implemented as finite state machines (FSM) in Verilog. Here we specify two finite state machines: one for generating trigger, and a second for controlling $AR_{en}$.
@@ 6.1. Sub-module FSMs
Fig. 9 shows the FSM for modifying AR via its corresponding bitvector $AR_{en}$. Each state corresponds to the status of $AR_{en}$:
@@
    Exec: state corresponding to execution that does not require reinstating or revoking a task to/from AR, and thus $AR_{en}$ remains unchanged in this state;
@@
This FSM initializes in the Exec state. Only upon a trigger does a transition into Revoke occur, and a transition from Exec into Reinstate occur only occurs when a software update completes (i.e., $PC$ being equal to $SW_{exit}$). Both Revoke and Reinstate states only execute for one cycle to update AR accordingly before transitioning to Exec. The transition from Exec into Revoke corresponds to LTL 7, and the transition from Exec into Reinstate corresponds to LTL 9.
Exec Revoke Reinstate else triggertrigger $PC=SW_{exit}$
@@ 6.2. Formal Verification
We use NuSMV (nusmv) as a model checker to determine whether PAIR design adheres to the specified properties and statements from Sec. 5 are upheld. We use Verilog2SMV (irfan2016verilog2smv) to convert Verilog specifications into nuSMV statements. Then, we formulate the LTL specifications in nuSMV to check the model. We execute the model checking atop an Intel(R) Core(TM) Ultra 7 155H (3.80 GHz) with 32.0 GB RAM using Ubuntu-20.04 under Windows Subsystem for Linux 2 (WSL 2). The model checker executed for 43.98 seconds with peak memory usage of 1.38GB to complete the verification.
@@ 7.1. Prototype configuration
To implement PAIR prototype, we use the Xilinx Vivado tool-set (vivado) as a development platform. PAIR hardware is written in the Verilog hardware description language and implements each of PAIR sub-modules according to the logic defined in Sec. 6, and its software is implemented in C. Using Vivado, we synthesize PAIR hardware, simulate behavior, evaluate area/energy costs, and deploy a PAIR-equipped openMSP430 (openmsp430) on the Basys3 prototyping board (basys3) We configure openMSP430 with 16KB of PMEM and 12KB of DMEM.
A custom linker is used to ensure all code and data of interest (from Fig. 2) are placed into contiguous memory regions that are monitored by PAIR. For the RTOS in our evaluation, we use RIOT (riot) due to its open-source availability and support for MCUs within our scope. In theory, PAIR concepts can apply to any RTOS with an identifiable API for kill and yield procedures.
For the IM, we instantiate a task-aware shadow stack module along with an indirect-branch table (IBT) for each task in hardware. Both interface with MCU memory to store their data and monitor MCU signals as required. It generates a violation upon illegal return instructions according to the shadow stack module and illegal forward branch instructions according to the indirect branch tables. Additionally, it monitors tasks’ stacks to ensure they do not interfere with each other’s temporary variables.
@@ 7.2. Fixed Costs
PAIR hardware cost is accounted for in the Look-up tables (LUTs) and flip-flop registers (FFs) added to the openMSP430 core. We compare this cost to that of closely related works in Fig. 10.
PAIRGAROTAACFAVRASEDUCCA EILID (CASU) AION (Sancus 2.0) 08008001,6001{,}600TotalLUTsFFs
PAIR requires fewer resources than all closely related works that deploy security extensions atop openMSP430. Although PAIR has a different goal than each of these works, this shows that it is cost-effective for low-end devices and is in line with related works that target the same class of devices. The most closely related works are EILID (eilid) and AION (aion). EILID implements CFI for low-end MCUs atop a hybrid (software-hardware) root of trust CASU (de2022casu) for memory immutability and trusted software updates. We note that when including IM in the cost, PAIR + IM requires more hardware resources than EILID. However, we note that PAIR is concerned with providing availability and run-time integrity simultaneously, whereas EILID’s primary focus is run-time integrity (discussed further in Sec. 8). AION builds atop Sancus (Sancus17) hardware-based TEE, which is more expensive due to requiring full TEE features and incorporating a cryptographic engine.
We also evaluate PAIR performance overheads by considering additional energy consumption. To evaluate energy overheads, we use Vivado synthesis tools to estimate PAIR’s power consumption on the Basys3 FPGA board. We find the openMSP430 core without PAIR consumes 71 mW of static power. Performing the same analysis on PAIR-equipped openMSP430, we observe 76 mW of static power, showing that PAIR incurs only an additional 5 mW in static power consumption.
@@ 7.3. Performance costs
We evaluate the performance costs of PAIR by evaluating the run-time and memory overheads when executing software atop PAIR. PAIR hardware does not impose any run-time overheads, since it is monitoring MCU signals without affecting the MCU’s critical path. The memory cost incurred by PAIR varies per application. For instance, $\mathsf{D_{\textsf{{PAIR}}}}$ size depends on the number of independent tasks. Additionally, since the IM may have a memory footprint, additional memory overhead is incurred while using PAIR that is not due to PAIR. Our IM instance memory usage for shadow stacks depends on the maximum call depth among all tasks, and IBT size depends on the total number of indirect forward edge control flow transfers among all tasks. To effectively evaluate PAIR memory usage, we select a set of applications from RIOT test examples and BEEBs embedded systems benchmark suite (beebs). We select RIOT test examples that create multiple tasks, and BEEBS applications that meet one or more of the following criteria:
@@
    has one or more indirect forward edges (jump/call);
@@
    has high or very high suitability for MCUs (beebs).
Table 1 shows the resulting memory usage due to PAIR alongside our IM instantiation.
Table 1. Memory usage (bytes) for applications with multiple tasks (MT), recursion (R), indirect forward edge control flows (IF), and high or very high suitability for MCUs (SUI) (beebs).
@@
Among the selected applications, PAIR memory usage is minimal, an increase ranging from just 4 to 16 bytes (0.024% ot 0.097% of available DMEM). Additional memory usage is required due to the IM, but it is still minimal: an increase ranging from 8 to 42 bytes (0.048% to 2.28% of the available data memory). PAIR by itself has a minimal fixed memory cost ($2+2\times N$ bytes), but incurs additional variable costs due to the IM that is installed.
@@ 8. Related Work
Many works have explored CFI in embedded and real-time systems (cfi_rts_survey). InsectACIDE (insectacide) proposes a method for CFI for ARM Cortex-M MCUs leveraging debugging extensions (arm_coresight) to record control flow transfers without incurring run-time overhead. Then, recorded control flow transfers are checked for integrity violations during idle times. RECFISH (recfish) demonstrates an RTOS-compatible CFI method for ARM Cortex-R MCUs by implementing shadow stacks and function-label checks via software instrumentation. FastCFI (fastcfi) also uses ARM extensions, but instead uses an FPGA to store and check a CFG. Regardless of their specificities, PAIR differs from each in that it accounts for the post-violation phase, allowing non-violating tasks to continue their execution.
@@
EILID (eilid) proposes a hybrid method for low-end MCUs and also evaluates atop openMSP430. It uses code instrumentation to implement trampolines that jump to a hardware-protected CFI monitor. PAIR +IM incurs more hardware overhead, while EILID incurs more run-time overhead and an increase in code size. Similar to PAIR, EILID triggers an interrupt upon violations and also enables software updates. However, EILID’s interrupt jumps directly to the software update function. PAIR’s interrupt instead disables a violating task before returning to the next high-priority benign task. Software updates may occur independently upon receiving an authenticated message. Due to parallel goals, PAIR and EILID could coexist, using PAIR to provide post-violation availability/integrity while using EILID to obtain a low-cost IM.
@@
Several related works address availability guarantees under full software compromise in MCUs (garota; acfa; traces; aion). Among them is AION (aion), which builds upon Sancus TEE (Sancus17) to provide strong availability guarantees to a trusted scheduler in the presence of compromised software. Like AION, we also built our prototype using openMSP430 as an MCU and RIOT as an RTOS. AION ensures the availability of a trusted Sancus-aware scheduler by incorporating hardware-based atomicity monitor and exception engine modules. While both AION and PAIR invoke the RTOS scheduler upon a violation signal, they pursue complementary objectives. AION aims to preserve the availability of a TEE-protected trusted scheduler to fairly allocate execution time to tasks, whereas PAIR focuses on preventing tasks that violate run-time integrity policies from executing until they are updated, thereby preserving availability to other non-violating tasks on the same platform. This is provided even if other software, including the scheduler, attempts to re-enter it. Given the orthogonal goals and commonality between MCU and RTOS platform for prototyping, PAIR and AION combined could provide a strong security service: AION’s trusted scheduler availability with PAIR’s enforcement against re-entering of violating tasks could return an architecture that ensures fair/timely scheduling and post-violation containment.
Active roots of trust (ARoTs) (acfa; traces; garota) ensure a contiguous region of memory (housing a trusted interrupt service routine) is guaranteed to execute upon its corresponding signal or upon MCU reset. Unlike ARoTs, which monitor a statically defined PMEM region, PAIR dynamically controls which tasks are granted availability.
@@ 9. Conclusion
In this work, we propose PAIR to preserve task availability and system-wide run-time integrity. By ensuring that only the offending task is prevented from executing, PAIR enables non-offending critical real-time tasks to continue executing. We designed, implemented, and evaluated a formally verified and open-source prototype of PAIR (repo), finding that its design has reasonable costs, incurring no runtime overhead while imposing minimal hardware and performance overheads.

### v2

@@ 7. Implementation & Verification
PAIR modules are implemented as two finite state machines (FSMs) controlling trigger=and $AR_{en}$, respectively.
@@ 7.1. Sub-module FSMs
Fig. 9 shows the FSM for modifying AR via the bitvector $AR_{en}$. Each state corresponds to the status of $AR_{en}$:
@@
    Monitor: state corresponding to execution that in which tasks are monitored. This state does not require reinstating or revoking a task to/from AR; thus, $AR_{en}$ remains unchanged in this state;
@@
This FSM initializes in the Monitor state. Only upon a trigger does a transition into Revoke occur, and a transition from Monitor into Reinstate occur only occurs when a software update completes (i.e., $PC$ being equal to $SW_{exit}$). Both Revoke and Reinstate states only execute for one cycle to update AR accordingly before transitioning to Monitor. The transition from Monitor into Revoke corresponds to LTL 7, and the transition from Monitor into Reinstate corresponds to LTL 9.
Monitor Revoke Reinstate else triggertrigger $PC=SW_{exit}$
@@ 7.2. Formal Verification
We use NuSMV (Cimatti et al., 2002) as a model checker to determine whether PAIR design adheres to the specified properties and statements from Sec. 6 are upheld. We use Verilog2SMV (Irfan et al., 2016) to convert Verilog specifications into nuSMV statements. Then, we formulate the LTL specifications in nuSMV to check the model. We execute the model checking atop an Intel(R) Core(TM) Ultra 7 155H (3.80 GHz) with 32.0 GB RAM using Ubuntu-20.04 under Windows Subsystem for Linux 2 (WSL 2). The model checker executed for 43.98 seconds with peak memory usage of 1.38GB to complete the verification.
@@ 8.1. Prototype configuration
To implement PAIR prototype, we use the Xilinx Vivado tool-set (Xilinx, 2017) as a development platform. PAIR hardware is written in the Verilog hardware description language and implements each of PAIR sub-modules according to the logic defined in Sec. 7, and its software is implemented in C. Using Vivado, we synthesize PAIR hardware, simulate behavior, evaluate area/energy costs, and deploy a PAIR-equipped openMSP430 (Girard, 2009) on the Basys3 prototyping board (Digilent, 2018) with 16KB of PMEM and 12KB of DMEM.
A custom linker is used to ensure all code and data of interest (from Fig. 2) are placed into contiguous memory regions that are monitored by PAIR. For the RTOS in our evaluation, we use RIOT (rio, 2010) due to its open-source availability, support for MCUs within our scope, and use in prior works (Alder et al., 2021). In theory, PAIR concepts can apply to any RTOS with an identifiable API for kill and yield procedures.
For the IM, we instantiate a task-aware shadow stack module along with an indirect-branch table (IBT) in hardware for CFI of return and jump instructions, respectively. Both interface with memory and MCU signals as required. Additionally, it monitors tasks’ stacks to ensure they do not interfere with each other’s temporary variables.
PAIRGAROTAACFAVRASEDUCCA EILID (CASU) AION (Sancus) 07507501,5001{,}500TotalLUTsFFs
@@ 8.2. Fixed Costs
PAIR hardware cost is accounted for in the Look-up tables (LUTs) and flip-flop registers (FFs) added to the openMSP430 core. We compare this cost to that of closely related works in Fig. 10. PAIR requires fewer resources than all closely related works that deploy security extensions atop openMSP430. Although PAIR has a different goal than each of these works, this shows that it is cost-effective for low-end devices and is in line with related works that target the same class of devices. The most closely related works are EILID (Jakkamsetti et al., 2025) and AION (Alder et al., 2021). EILID implements CFI for low-end MCUs atop a hybrid (software-hardware) root of trust CASU (De Oliveira Nunes et al., 2022) for memory immutability and trusted software updates. We note that when including IM in the cost, PAIR + IM requires more hardware resources than EILID. However, we note that PAIR is concerned with providing availability and run-time integrity simultaneously, whereas EILID’s primary focus is run-time integrity (discussed further in Sec. 9). AION builds atop Sancus (Noorman et al., 2017) hardware-based TEE, which requires a more extensive set of security features.
We also evaluate PAIR performance overheads by considering additional energy consumption. To evaluate energy overheads, we use Vivado synthesis tools to estimate PAIR’s power consumption on the Basys3 FPGA board. We find the openMSP430 core without PAIR consumes 71 mW of static power and 76 mW with PAIR, showing that PAIR incurs only an additional 5 mW in static power consumption.
@@ 8.3. Performance costs
We evaluate the performance costs of PAIR using the run-time and memory overheads when executing software atop PAIR. PAIR hardware does not impose any run-time overheads, since it is monitoring MCU signals without affecting the MCU’s critical path. The memory cost incurred by PAIR varies per application. For instance, $\mathsf{D_{\textsf{{PAIR}}}}$ size depends on the number of independent tasks. Memory cost may also vary based on the particular IM in use. Our IM instance memory usage for shadow stacks depends on the maximum call depth among all tasks, and IBT size depends on the total number of indirect forward edge control flow transfers among all tasks. To effectively evaluate PAIR memory usage, we select a set of applications from RIOT test examples and BEEBs embedded systems benchmark suite (Pallister et al., 2019). We select RIOT test examples that create multiple tasks, and BEEBS applications that meet one or more of the following criteria:
@@
    has one or more indirect forward branches (jump/call);
@@
    has high or very high suitability for MCUs (Pallister et al., 2019).
Table 1 shows the resulting memory usage.
Table 1. Memory usage for applications with multiple tasks (MT), recursion (R), indirect forward branches (IF), and high/very high MCU suitability (SUI) (Pallister et al., 2019).
@@
Among the selected applications, PAIR memory usage is minimal, an increase ranging from just 4 to 16 bytes (0.024% to 0.097% of available DMEM). Additional memory usage is required due to the IM, but it is still minimal: an increase ranging from 8 to 42 bytes (0.048% to 2.28% of the available data memory). PAIR by itself has a minimal fixed memory cost ($2+2\times N$ bytes), but incurs additional variable costs due to the IM that is installed.
[…]
@@ 9. Related Work
Many works have explored CFI in embedded and real-time systems (Mishra et al., 2022). InsectACIDE (Wang et al., 2024) proposes a method for CFI for ARM Cortex-M MCUs leveraging CoreSight tracing extensions to record control flow transfers without incurring run-time overhead. Then, recorded control flow transfers are checked for integrity violations during idle times. RECFISH (Walls et al., 2019) demonstrates an RTOS-compatible CFI method for ARM Cortex-R MCUs through shadow stacks and function-label checks via software instrumentation. FastCFI (Feng et al., 2021) also uses ARM extensions, but instead uses an FPGA to store and check a CFG. Regardless of their specificities, PAIR differs from each in that it accounts for the post-violation phase, allowing non-violating tasks to continue their execution.
@@
EILID (Jakkamsetti et al., 2025) proposes a hybrid method for low-end MCUs and also evaluates atop openMSP430. It uses code instrumentation to implement trampolines that jump to a hardware-protected CFI monitor. PAIR +IM incurs more hardware overhead, while EILID incurs more run-time overhead and an increase in code size. Similar to PAIR, EILID triggers an interrupt upon violations and also enables software updates. However, EILID’s interrupt jumps directly to the software update function. PAIR’s interrupt instead disables a violating task before returning to the next high-priority benign task. Due to parallel goals, PAIR and EILID could coexist, using PAIR to provide post-violation availability/integrity with EILID as a low-cost IM.
Active roots of trust (ARoTs) (Caulfield et al., 2023, 2024; Aliaj et al., 2022; Alder et al., 2021; Gülmez et al., 2023; Lengert et al., 2026) ensure a trusted function is guaranteed to execute upon a corresponding hardware signal or upon MCU reset. The concept of ARoTs was proposed in GAROTA (Aliaj et al., 2022), a formally-verified architecture that provides two properties regarding the execution of a trusted interrupt service routine (ISR): (1) the guaranteed triggering of the trusted ISR upon its interrupt source (e.g., GPIO, timer), and (2) the guaranteed re-triggering of the trusted ISR upon boot or reset. Several works built upon GAROTA to realize ARoTs for various purposes (Caulfield et al., 2023; Lengert et al., 2026; Caulfield et al., 2024).
@@
AION (Alder et al., 2021) is an ARoT builds upon Sancus TEE (Noorman et al., 2017) to provide strong availability guarantees to a trusted scheduler in the presence of compromised software. Like AION, we also built our prototype using openMSP430 as an MCU and RIOT as an RTOS. AION ensures the availability of a trusted Sancus-aware scheduler by incorporating hardware-based atomicity monitor and exception engine modules. While both AION and PAIR invoke the RTOS scheduler upon a violation signal, they pursue complementary objectives. AION aims to preserve the availability of a TEE-protected trusted scheduler to fairly allocate execution time to tasks, whereas PAIR focuses on preventing tasks that violate run-time integrity policies from executing until they are updated, thereby preserving availability to other non-violating tasks on the same platform. This is provided even if other software, including the scheduler, attempts to re-enter it. Given the orthogonal goals, PAIR and AION combined could provide a strong security service: AION’s trusted scheduler with PAIR’s enforcement against re-entering of violating tasks could return both fair scheduling and post-violation containment.
SDRaD (Gülmez et al., 2023) leverages Intel PKU to compartmentalize applications and to trigger rollback of execution state upon a violation. SDRaD is tailored towards restoring prior execution context with custom handlers to avoid replay of the attack and targets rich application software on COTS processors, whereas PAIR targets resource-constrained MCUs operating under hard real-time constraints, enforcing post-violation task containment entirely in hardware, and explicit integration with RTOS scheduling semantics.
10. Discussion & Future Work
Side-channel attacks. PAIR’s software update routine authenticates update requests using a secret key. A naive implementation may be vulnerable to timing side-channel attacks if it enables any secret-dependent execution paths. This can be mitigated by ensuring the authentication routine is implemented as constant-time code. Cache-timing side-channels are a common concern for cryptographic routines on general-purpose processors. However, this is not applicable to devices in PAIR’s scope, which often do not incorporate caches. However, this is an important consideration when extending to more complex processors in future work.
Multi-Core Embedded Systems. Extending PAIR to multi-core embedded systems introduces challenges. For example, AR state must be maintained atomically across concurrent contexts, and kill-and-yield semantics must account for cross-core scheduling. An extension of PAIR to support multi-core is an intriguing direction for future work.
Interdependent Tasks. PAIR assumes tasks are functionally independent. It is possible that an application is designed such that its tasks coordinate via locks, mutexes, or semaphores. As such, removing a violating task from AR mid-execution may leave shared locks unreleased, causing deadlock. Addressing this requires further consideration of RTOS locking primitives, including monitoring of its access and ability to unlock in special cases from PAIR’s SW protected region. We leave this as a direction for future work.
Task Criticality. PAIR provides an effective balance for scenarios in which preventing a task from continuing is safer than allowing it to continue in an unexpected state, regardless of the task’s criticality. Although applicable to many scenarios (as described in Section 3), there are scenarios in which preventing a compromised high-criticality task from continuing may still be unideal (e.g., automotive vehicles). Future work could explore enhancing PAIR SW to differentiate violation handling based on criticality: lower-criticality tasks are immediately removed from AR, while higher-criticality tasks are re-executed once before removal, leveraging aspects of PAIR and related works (Alder et al., 2021; Gülmez et al., 2023; Aliaj et al., 2022).
@@ 11. Conclusion
In this work, we propose PAIR to preserve task availability and system-wide run-time integrity. By ensuring that only the offending task is prevented from executing, PAIR enables non-offending critical real-time tasks to continue executing. We designed, implemented, and evaluated a formally verified and open-source prototype of PAIR  [^1] To be made public after peer review, finding that its design has reasonable costs, incurring no runtime overhead while imposing minimal hardware and performance overheads.